11. Find the area of the shaded region under the standard normal curve. If​ convenient, use technology to find the area. z -2.13 0 A normal curve is over a horizontal z-axis and is centered at 0. Vertical line segments extend from the horizontal axis to the curve at negative 2.13 and 0. The area under the curve between negative 2.13 and 0 is shaded. The area of the shaded region is nothing. ​(Round to four decimal places as​ needed.)
Answer #1

11)

Solution :

Using standard normal table,

P(-2.13 < z < 0) = P(z < 0) - P(z < -2.13) = 0.5 - 0.0166 = 0.4834

The area of the shaded region is 0.4834